# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v3

@@ IV-C Comparison between the Sliding DFT-based recovery and HoD-based recovery
We now compare the performance of the proposed sliding DFT recovery against an existing unfolding method for modulo sampling. Note that it is not practical to unfold $f(t)$ in (22) using the DFT-based methods described in [10, 13, 14, 11, 36, 48] due to its large $N_{0}$ and its lack of periodicity. Moreover, the thresholding-based recovery method in [28, Algorithm 1] is unsuitable for the given $f(t)$, as it requires a minimum separation between folding times to function correctly. The low-sampling-rate variant [28, Algorithm 2] is likewise impractical due to its $\mathcal{O}(N_{0}^{2})$ computational complexity. As a benchmark, we compare our proposed sliding DFT-based recovery method against the HoD modulo recovery approach [6, 7].
@@
Figure 9: Simulated MSE vs. oversampling factor of the sliding DFT-based recovery and HoD-based recovery for $b=3,4,5$.
@@
Figure 9 shows the simulated MSE performance of the proposed sliding DFT-based recovery method compared to the HoD-based recovery for different values of $\mathrm{OF}$ and $b$. The parameters for the sliding DFT-based method are the same as in Section IV-B, except with $\delta_{\mathrm{SL}}=\frac{\pi}{32}$. For both methods, the modulo threshold $\lambda^{\prime}$ is set according to (18) and the test signal $f(t)$ is given in (22). When $b=4$ or $b=5$, the proposed method successfully unfolds the modulo ADC output across all $\mathrm{OF}$ values considered. In contrast, the HoD-based recovery shows very high MSE for $4<\mathrm{OF}<15$ when $b=4$, and for $4<\mathrm{OF}<8$ when $b=5$. These results highlight the advantage of the proposed sliding DFT-based recovery over HoD-based recovery in the low-sampling-rate, low-resolution regime. The performance advantage of our proposed sliding DFT-based recovery over HoD could be attributed to the availability of the 1-bit folding information.
@@
We also note that both methods fail to unfold the modulo ADC output when $b=3$, suggesting that $b>3$ may be a necessary condition for the successful operation of the sliding DFT-based recovery in its current form. Further research is needed to confirm this conjecture. Alternatively, employing other quantizer architectures in conjunction with modulo sampling may enable sliding DFT-based unfolding at resolutions below $b=4$. For instance, the one-bit unlimited sampling approach in [55] uses a 1-bit sigma–delta quantizer together with modulo sampling to implement a 1-bit modulo ADC.
@@ V Summary and Future Directions
In this work, we considered a modulo ADC system in which the output samples are associated with a 1-bit folding information signal $c[n]$. The unfolded samples are obtained using the sliding DFT-based recovery scheme described in Section II-C. The advantage of our proposed unfolding scheme over existing Fourier-based recovery algorithms for modulo sampling is the significant reduction in the length of the observation window. This reduction in the observation time has a positive impact in the algorithm’s computational complexity. We also provide sufficient conditions on the oversampling factor and ADC resolution to attain a certain MSE performance guarantee. When these sufficient conditions are met, the MSE performance of the modulo ADC which uses our proposed recovery method is better compared to that of conventional ADCs. Moreover, we also showed that spectral leakage affects the asymptotic MSE growth rate of our sliding DFT recovery method. These theoretical results are substantiated by the numerical experiments conducted in Section IV. Under the same modulo threshold setting and $b=4,5$, the proposed sliding DFT recovery method also outperformed the HoD-based recovery at low values of the oversampling factor.
It is worth mentioning that while the proposed unfolding scheme relies on the availability of 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be done by performing robust sparse recovery of $\underline{\tilde{z}^{(i)}_{w}}[n]$ using a sliding DFT version of the LASSO-$B^{2}R^{2}$ [13] or a sliding DFT version of the OMP-based unfolding in [48]. This idea, together with its performance guarantees, is considered a potential subject of future work. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution are not tight. Hence, another future direction is to tighten these sufficient conditions. The modulo sampling architecture considered in this work did not take into account non-ideal modulo nonlinearities such as hysteresis and folding transients. Thus, we also intend to apply the sliding DFT recovery to generalized modulo sampling models and analyze how modulo hysteresis and folding transients affect the algorithm’s performance. Lastly, it is also of interest to apply this proposed recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers, radars, etc.).
@@ Appendix B Proof of Proposition 2
for all $i$. The first line follows from multiplying the number of level crossings in a single level and the number of folding levels. The second line comes from the trivial lower bound of the floor function, i.e., $\lfloor x\rfloor\leq x$. Since the bound in equation (B) is independent of $i$, this bound is also an upper bound for $\underset{i}{\max}\;|\mathcal{S}_{i}|$. Using equation (B), the following inequality holds:
@@
$\mathrm{OF}\geq\frac{N}{N-\frac{N}{\mathrm{OF}}\left(1+\frac{\|f(t)\|_{\infty}}{\lambda^{\prime}}\right)-K_{\mathrm{SL}}}$
@@
as the modulo threshold satisfies the inequality in (27). However, since $\lambda^{\prime}\leq\|f(t)\|_{\infty}$, then
$\mathrm{OF}\cdot\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2\geq 1.$
@@ Appendix C Proof of Theorem 1
$=$ $\;\left(1+\|\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}^{c}}\|_{\infty}\right)\cdot\left(\frac{6\lambda^{\prime}}{2^{b}-2}\right).$
The first line comes from the fact that $\underline{\tilde{z}}^{(i)}[n]=\underline{{z}}^{(i)}[n]=0$ for $n\notin\mathcal{S}_{i}$. The second and third lines come from equations (6) and (7). Because the last $\frac{\alpha N}{2}$ samples of $\tilde{z}_{w}^{(i-1)}[n]$ is added to the first $\frac{\alpha N}{2}$ samples of $\tilde{z}_{w}^{(i)}[n]$, the effect of the window function is removed. Consequently, the subscript $w$ is dropped in the second line. The fourth line holds because $\mathbf{V}_{\mathcal{S}_{i}}$ is full column rank by Proposition 2 and $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}\underline{\boldsymbol{\epsilon}_{\mathcal{S}_{i}}}+\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}^{c}}\underline{\boldsymbol{\epsilon}_{\mathcal{S}_{i}^{c}}}=\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}\underline{\boldsymbol{\epsilon}}$. The fifth line follows from the property $\|\mathbf{A}\mathbf{x}\|_{\infty}\leq\|\mathbf{A}\|_{\infty}\|\mathbf{x}\|_{\infty}$. The sixth line holds because the triangle dither induces a quantization noise $\epsilon[n]$ whose range of amplitude is $\left(-\frac{3\lambda}{2^{b}},+\frac{3\lambda}{2^{b}}\right)$. Consequently, $\underline{\epsilon}[n]$ has amplitude in $\left(-\frac{6\lambda^{\prime}}{2^{b}},+\frac{6\lambda^{\prime}}{2^{b}}\right)$. Finally, the last line comes from the relationship between the modulo threshold $\lambda^{\prime}$ and the ADC dynamic range parameter $\lambda$.
@@
$\frac{6\lambda^{\prime}}{2^{b}-2}(1+M)<\lambda^{\prime},$
@@
$=$ $\frac{(\lambda^{\prime})^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}}\left(1+\frac{\delta_{\mathrm{SL}}\cdot\mathrm{OF}}{\pi}\right).$

### v4

@@ IV-C Comparison between the STFT-based recovery and HoD-based recovery
We now compare the performance of the proposed STFT-based recovery method with an existing unfolding method for modulo sampling. Note that it is not practical to unfold $f(t)$ in (29) using the DFT-based methods described in [6, 45, 4, 1, 46, 5] because of its large $N_{0}$ and its lack of periodicity. Moreover, the thresholding-based recovery method in [19, Algorithm 1] is unsuitable for the given $f(t)$, because it requires a minimum separation between folding times to function correctly. The low-sampling rate variant [19, Algorithm 2] is likewise impractical because to its $\mathcal{O}(N_{0}^{2})$ computational complexity. As a benchmark, we compared our proposed STFT-based recovery method with the HoD modulo recovery approach [7, 9].
@@
Figure 11: Simulated MSE vs. oversampling factor of the STFT-based recovery for $b=3,4,5$. Superimposed in the figure are the simulated MSE curves of the HoD-based recovery.
@@
Figure 11 shows the simulated MSE performance of the proposed STFT-based recovery method compared with the HoD-based recovery method for different values of $\mathrm{OF}$ and $b$. The parameters for the STFT-based method are the same as those in Section IV-B, except for $\delta_{\mathrm{SL}}=\frac{3\pi}{64}$. For both methods, the modulo threshold $\lambda^{\prime}$ is set according to (23), and the test signal $f(t)$ is given by (29). When $b=4$ or $b=5$, the proposed method successfully unfolded the modulo ADC output across all $\mathrm{OF}$ values considered. In contrast, the HoD-based recovery shows very high MSE values for $\mathrm{OF}<14$ when $b=4$, and for $\mathrm{OF}<7$ when $b=5$. These results highlight the advantage of the proposed STFT-based recovery method over the HoD-based recovery in the low-sampling rate, low-resolution regime. The performance advantage of our proposed STFT-based recovery method over the HoD-based recovery method could be attributed to the availability of 1-bit folding information.
@@
We also note that both methods fail to unfold the modulo ADC output when $b=3$, suggesting that $b>3$ may be a necessary condition for the successful operation of STFT-based recovery in its current form. Further research is needed to confirm this hypothesis. Alternatively, employing other quantizer architectures in conjunction with modulo sampling may enable STFT-based unfolding at resolutions below $b=4$. For instance, the one-bit unlimited sampling approach in [24, 40] uses a 1-bit sigma–delta quantizer together with modulo sampling to implement a 1-bit modulo ADC.
@@ V Summary and Future Directions
In this work, we considered a modulo ADC system in which the output samples are associated with a 1-bit folding information signal $c[n]$. Unfolded samples were obtained using the STFT-based recovery scheme described in Section II-C. The advantage of our proposed recovery scheme over existing Fourier-based recovery algorithms for modulo sampling is the significant reduction in the observation window length. This reduction in observation time reduces the computational complexity of the algorithm. We also provided sufficient conditions for the oversampling factor and ADC resolution to guarantee a certain MSE performance. When these sufficient conditions are met, the MSE performance of the modulo ADC that uses our proposed recovery method is better than that of the conventional ADCs. Moreover, we demonstrated that spectral leakage affects the asymptotic MSE growth rate of our STFT-based recovery method. These theoretical results are substantiated by the numerical experiments conducted in Section IV. Under the same modulo threshold setting and $b=4,5$, the proposed STFT-based recovery method outperforms the HoD-based recovery method at low oversampling factor values. These findings help connect classical sampling theory with modulo quantization frameworks and point to new directions for designing low-resolution front ends with established performance guarantees.
It is worth mentioning that while the proposed unfolding scheme relies on the availability of a 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be achieved by performing robust sparse recovery of $\underline{\tilde{z}^{(i)}_{w}}[n]$ using an STFT-based version of LASSO-$B^{2}R^{2}$ [45] or an STFT-based version of OMP-based unfolding in [5]. This idea, together with its performance guarantees, is considered a potential subject for future research. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution were not tight. Hence, another future direction is to tighten these conditions.
It is important to note that the performance guarantees presented in this study assume ideal modulo nonlinearities, that is, instantaneous and perfectly accurate folding without hysteresis or folding transients. This assumption represents a limitation of the current analysis, because real hardware implementations may deviate from this ideal behavior. Thus, we also intend to apply STFT-based recovery to generalized modulo sampling models and analyze how modulo hysteresis and folding transients affect the performance of the algorithm. Another direction is experimental validation of the proposed method under realistic analog impairments. Practical ADCs experience thermal noise, front-end nonlinearities, temperature drift, and other impairments that are not captured by the present theoretical model. Finally, it is of interest to apply this recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers and radars).
@@ Appendix B Proof of Proposition 2
for all $i$. The first line follows from multiplying the number of level crossings in a single level and the number of folding levels. The second line comes from the trivial upper bound of the floor function, i.e., $\lfloor x\rfloor\leq x$. Since the bound in equation (B) is independent of $i$, this bound is also an upper bound for $\underset{i}{\max}\;|\mathcal{S}_{i}|$. Using equation (B), the following inequality holds:
@@
$\mathrm{OF}\geq\frac{N}{N-\frac{N}{\mathrm{OF}}\left(1+\frac{\|f(t)\|_{\infty}}{\lambda^{\prime}}\right)-K_{\mathrm{SL}}}$ (34)
@@
as the modulo threshold satisfies the inequality in (35). However, since $\lambda^{\prime}\leq\|f(t)\|_{\infty}$, we get
$\mathrm{OF}\cdot\left[1-\frac{K_{\mathrm{SL}}}{N}\right]-2\geq 1.$ (36)
@@ Appendix C Proof of Theorem 1
$=$ $\;\left(1+\|\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}^{c}}\|_{\infty}\right)\cdot\left(\frac{6\lambda^{\prime}}{2^{b}-2}\right).$ (37)
The first line comes from the fact that $\underline{\tilde{z}}^{(i)}[n]=\underline{{z}}^{(i)}[n]=0$ for $n\notin\mathcal{S}_{i}$. The second and third lines come from equations (8) and (9). Because the last $\frac{\alpha N}{2}$ samples of $\tilde{z}_{w}^{(i-1)}[n]$ are added to the first $\frac{\alpha N}{2}$ samples of $\tilde{z}_{w}^{(i)}[n]$, the effect of the window function is removed. Consequently, the subscript $w$ is dropped in the second line. The fourth line holds because $\mathbf{V}_{\mathcal{S}_{i}}$ is full column rank by Proposition 2 and $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}}\underline{\boldsymbol{\epsilon}_{\mathcal{S}_{i}}}+\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}_{\mathcal{S}_{i}^{c}}\underline{\boldsymbol{\epsilon}_{\mathcal{S}_{i}^{c}}}=\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\mathbf{V}\underline{\boldsymbol{\epsilon}}$. The fifth line follows from the property $\|\mathbf{A}\mathbf{x}\|_{\infty}\leq\|\mathbf{A}\|_{\infty}\|\mathbf{x}\|_{\infty}$. The sixth line holds because the triangle dither induces a quantization noise $\epsilon[n]$ whose range of amplitude is $\left(-\frac{3\lambda}{2^{b}},+\frac{3\lambda}{2^{b}}\right)$. Consequently, $\underline{\epsilon}[n]$ has amplitude in $\left(-\frac{6\lambda^{\prime}}{2^{b}},+\frac{6\lambda^{\prime}}{2^{b}}\right)$. Finally, the last line comes from the relationship between the modulo threshold $\lambda^{\prime}$ and the ADC dynamic range parameter $\lambda$.
@@
$\frac{6\lambda^{\prime}}{2^{b}-2}(1+M)<\lambda^{\prime},$ (38)
@@
$=$ $\frac{(\lambda^{\prime})^{2}}{\mathrm{OF}\cdot(2^{b}-2)^{2}}\left(1+\frac{\delta_{\mathrm{SL}}\cdot\mathrm{OF}}{\pi}\right).$ (39)